OpenStudy (anonymous):

Solve by extracting roots. Assume both a and b are positive (ax+b)2=81

OpenStudy (campbell_st):

start with taking the square root of both sides then subtract b then divide both sides by a it will look a little like the general quadratic formula.

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

i didnt gwt the answer @campbell_st

OpenStudy (agent0smith):

\[(ax+b)^2=81\]take square root of both sides \[ax+b=\pm 9\] Now try to finish by subtracting b from both sides, then dividing by a.

OpenStudy (anonymous):

\[(ax+b)^{2}=81\]\[ax+b=\pm9\] Case 1 \[ax+b=9 => ax=9-b\]\[x=\frac{ 9-b }{ a }\] Case 2 \[ax+b=-9 => ax=-9-b\] \[\frac{ -9-b }{ a}\]